A bottle contains 255 coins
1/3 of the coins are £1 coins
110 of the coins are 50p coins
the rest of the coins are 20p coins
work out the total value of the coins contained in the bottle
1
Expert's answer
2018-10-02T10:58:08-0400
Number of £1 coins: N_1=255/3=85.
Number of 50p coins: N_2=110.
Number of 20p coins: N_3=255-85-110=60.
The total value of the coins:
T=N_1*1+N_2*0.5+N_3*0.2=85*1+110*0.5+60*0.2=£152.